OpenStudy (shamil98):

Use power series to solve for y'' + y = 0.

OpenStudy (kc_kennylau):

"Use power series to..." What if no.

OpenStudy (kc_kennylau):

xD

OpenStudy (shamil98):

It's quite long, and i'm actually pondering on how it's done... It's the next section after nonhomogenous equations ( which I have yet to learn to xD)

OpenStudy (kainui):

\[y=\sum_{n=0}^{\infty}a_nx^n\] \[y''=\sum_{n=0}^{\infty}(n+1)(n+2)a_{n+2}x^n\] \[y''+y=0=\sum_{n=0}^{\infty}x^n[a_n+(n+1)(n+2)a_{n+2}]\] since x is variable, that part can't be assumed to be 0. So we have to assume the sum of the other stuff is 0. \[0=a_n+(n+1)(n+2)a_{n+2}\] This gives you a recursion relation which amounts to finding even and odds which are your two constants for your fundamental solutions which are determined by a_0 and a_1.

OpenStudy (kc_kennylau):

wait, isn't y'' [SUM] n(n-1)x^(n-2)?

OpenStudy (kc_kennylau):

http://tutorial.math.lamar.edu/Classes/DE/SeriesSolutions.aspx Just googled this and learnt it by myself xD (I always do this when I meet new terms)

OpenStudy (kainui):

See, when you take the derivative of a power series, just increase the index from n=0 to n=1 since you know your n=0 term will go to zero since it's a constant term. That way when you get to the next thing you have the same index.

OpenStudy (kc_kennylau):

Oh I see I get it

OpenStudy (kainui):

\[y=\sum_{n=0}^{\infty}a_nx^n=a_0+\sum_{n=1}^{\infty}a_{n+1}x^{n+1}\]

OpenStudy (kc_kennylau):

@kainui Thanks for teaching me :D

OpenStudy (kainui):

No problem, that's why I come here. =P

OpenStudy (kc_kennylau):

but @shamil98 is gone :/

OpenStudy (shamil98):

Oh, I get it, over thought it a bit at first.. (sleep-deprivation kickin in xD)

OpenStudy (shamil98):

What is the next step to solve for it? or is that it?

OpenStudy (kc_kennylau):

I think @Kainui means that there has to be more information

OpenStudy (kc_kennylau):

I'm new in this matter

OpenStudy (anonymous):

a good practice...\[y''+xy'+(x^2+2)y=0 \] around \[x_0=0\]

OpenStudy (kc_kennylau):

Using whatever approach? @Jonask

OpenStudy (anonymous):

yes...it was explained above very well

OpenStudy (kainui):

So yeah, we are definitely not done, we need to find a recursion relation. I just got lazy cause it's x-mas and I figured that's pretty much the easy part. \[0=a_n+(n+2)(n+1)a_{n+2}\] \[-a_n/(n+2)(n+1)=a_{n+2}\] So now you plug in basically n=0 to find the recursion relation for the even terms, n=1 for the odd. Notice that we're going to obviously have a (-1)^n since it oscillates between positive and negative terms. Since each one up we go, we'll be continually multiplying (n+1)(n+2)(n+3)(n+4) etc... the bottom part will be a factorial. So then we can pretty much say after some mathmagic... (actually I just wrote in the power series for sine and cosine from memory, so they might be slightly wrong.) \[y=a_0\sum_{n=0}^{\infty}\frac{ (-1)^nx^{2n} }{ (2n)! }+a_1\sum_{n=0}^{\infty}\frac{ (-1)^nx^{2n+1} }{ (2n+1)! }\]

OpenStudy (shamil98):

thanks! I think i understood it. (i'll re-check this when I do have some sleep.)

OpenStudy (solomonzelman):

you forgot the 1st quotation mark.

OpenStudy (solomonzelman):

\["why"+why=0\]This?

OpenStudy (shamil98):

y'' means the second derivative , solomon.

OpenStudy (kc_kennylau):

nope it's y differentiated twice

OpenStudy (kc_kennylau):

@SolomonZelman

OpenStudy (solomonzelman):

Yeah I looked it up before typing....

OpenStudy (kainui):

The good thing about power series solutions to differential equations is that you can easily make up your own problems for yourself to solve. Obviously the answer to this one is sinx and cosx since their second derivatives are just the negative versions of themselves, but as long as its any differential equation you've found the answer to by some other method you should be fine if you're practicing the method.